The letters of the word "OKOTOKS" are arranged. The number of arrangements that begins and ends with a vowel is ?
Answer by Theo(13342)

if you put the vowels on the end, the number of ways you can do that is 1 because all the vowels are the same.

if you take away two of the o's because they're on the ends, then you are left with ktoks.

these can be arranged in 5! / 2! ways = (5*4*3*2!)/(2!) = 5*4*3 = 60 ways

60 ways in the middle * 1 way on the ends = 60 total ways.

that's my take.